OpenStudy (anonymous):

What is the domain of F(x)= x over 2x^2-5x-3

OpenStudy (ipwnbunnies):

If you plug in any x value, will you get an answer?

OpenStudy (ipwnbunnies):

In other words, is there any x-value or values where the function is undefined?

OpenStudy (anonymous):

well the answers are x does not =0 x does not =-1/2, x does not =3 x does not =-3, x does not =1/2

OpenStudy (ipwnbunnies):

What. o.o

OpenStudy (ipwnbunnies):

Oh, I read that problem wrong.

OpenStudy (ipwnbunnies):

We want to find the x values where the function is undefined. We can figure that out by setting the denominator equal to 0. Remember, anything divided by 0 is undefined.

OpenStudy (anonymous):



OpenStudy (ipwnbunnies):

No. There is a denominator in the function. The domain will not include the x-values that make the denominator 0.

OpenStudy (anonymous):

oh that was just one of the answers....like is this a problem where you have to factor things out and then see what x can't equal

OpenStudy (ipwnbunnies):

Yeah, you would have to factor the denominator, solve for x. Those x values won't be included in the domain.

OpenStudy (anonymous):

Yeah that's what I thougt...thanks